# Elliptic Quantum Toroidal Algebra $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor}),$ Vertex Operators and $L$-operators Hitoshi Konno and Andrey Smirnov

### v1

@@ 4.1 OPE of the elliptic currents
$<x^{-}(u)x^{-}(v)>=\frac{(v/u,\hbar v/u,pv/t_{1}u,pv/t_{2}u;p)_{\infty}}{(t_{2}v/u,t_{1}v/u,pv/\hbar u,pv/u;p)_{\infty}}.$
@@
$<x^{+}(u)x^{+}(v)>^{sym}=t_{1}^{-1}\frac{(p^{*}t_{2}v/u,p^{*}t_{2}u/v,v/u,u/v;p^{*})_{\infty}}{(p^{*}\hbar v/u,p^{*}\hbar u/v,v/t_{1}u,u/t_{1}v;p^{*})_{\infty}},$ (4.3)
$<x^{-}(u)x^{-}(v)>^{sym}=t_{1}^{-1}\frac{(pt_{1}^{-1}v/u,pt_{1}^{-1}u/v,\hbar v/u,\hbar u/v;p)_{\infty}}{(pv/u,pu/v,t_{2}v/u,t_{2}u/v;p)_{\infty}},$ (4.4)
@@
$<x^{+}(u)x^{+}(v)>^{sym}_{0}=\frac{(v/u,u/v,v/\hbar u,u/\hbar v;p^{*})_{\infty}}{(v/t_{1}u,u/t_{1}v,v/t_{2}u,u/t_{2}v;p^{*})_{\infty}},$ (4.7)
$<x^{-}(u)x^{-}(v)>^{sym}_{0}=\frac{(v/u,u/v,\hbar v/u,\hbar u/v;p)_{\infty}}{(t_{1}v/u,t_{1}u/v,t_{2}v/u,t_{2}u/v;p)_{\infty}},$ (4.8)
$<x^{+}(u)x^{+}(v)>^{non-sym}_{0}=\frac{\theta^{*}(t_{2}v/u)\theta^{*}(t_{1}v/u)}{\theta^{*}(v/u)\theta^{*}(\hbar v/u)},$ (4.9)
$<x^{-}(u)x^{-}(v)>^{non-sym}_{0}=\frac{\theta(t_{1}u/v)\theta(t_{2}u/v)}{\theta(u/v)\theta(\hbar u/v)}.$ (4.10)
@@
4.2 Vertex operators for ${\cal M}(n,r)$
We start from the vertex operators corresponding to the case $r=1$ i.e. $\mathrm{Hilb}^{n}({\mathbb{C}}^{2})$. Let us define
$\Phi_{\emptyset}(u):=(-u)^{-\alpha}e^{-\Lambda_{0}^{\perp}}\;\widetilde{\Phi}_{\emptyset}(u),$ (4.11)
$\Psi^{*}_{\emptyset}(u):=(-u)^{\alpha}e^{\Lambda_{0}^{\perp}-Q}\;\widetilde{\Psi}^{*}_{\emptyset}(u),$ (4.12)
@@ 4.2 Vertex operators for ${\cal M}(n,r)$
$\widetilde{\Phi}_{\emptyset}(u)=\exp\left\{-\sum_{m>0}\frac{1}{\kappa_{m}}\alpha^{\prime}_{-m}(\hbar^{1/2}u)^{m}\right\}\exp\left\{\sum_{m>0}\frac{1}{\kappa_{m}}\alpha^{\prime}_{m}(\hbar^{1/2}u)^{-m}\right\},$ (4.13)
$\widetilde{\Psi}^{*}_{\emptyset}(u)=\exp\left\{\sum_{m>0}\frac{1}{\kappa_{m}}\alpha_{-m}(\hbar^{1/2}u)^{m}\right\}\exp\left\{-\sum_{m>0}\frac{1}{\kappa_{m}}\alpha_{m}(\hbar^{1/2}u)^{-m}\right\}.$ (4.14)
@@
$\Psi^{*}_{\emptyset}(u)x^{+}(v)=-\frac{\theta^{*}(v/u)}{\theta^{*}(\hbar u/v)}x^{+}(v)\Psi^{*}_{\emptyset}(u),$ (4.15)
$\Phi_{\emptyset}(u)x^{-}(v)=-\frac{\theta(\hbar u/v)}{\theta(v/u)}x^{-}(v)\Phi_{\emptyset}(u),$ (4.16)
$x^{+}(v)\Phi_{\emptyset}(u)=\Phi_{\emptyset}(u)x^{+}(v),\qquad x^{-}(v)\Psi^{*}_{\emptyset}(u)=\Psi^{*}_{\emptyset}(u)x^{-}(v),$ (4.17)
$\mathfrak{z}\Phi_{\emptyset}(u)=\hbar^{-1}\Phi_{\emptyset}(u)\mathfrak{z},\qquad\Psi^{*}_{\emptyset}(u)\mathfrak{z}^{*}=\hbar\;\mathfrak{z}^{*}\Psi^{*}_{\emptyset}(u),$ (4.18)
$[x^{\pm}(x),\mathfrak{z}]=[\Psi^{*}_{\emptyset},\mathfrak{z}]=0,\qquad[x^{\pm}(x),\mathfrak{z}^{*}]=[\Phi_{\emptyset}(u),\mathfrak{z}^{*}]=0.$ (4.19)
@@
$\Phi(u)\ :\ {\cal F}^{(1,N)}_{v}\ \to\ {\cal F}^{(0,1)}_{u}\widetilde{\otimes}{\cal F}^{(1,N-1)}_{-v/u},$
$\Psi^{*}(u)\ :\ {\cal F}^{(1,N)}_{v}\widetilde{\otimes}{\cal F}^{(0,1)}_{u}\ \to\ {\cal F}^{(1,N+1)}_{-vu}$
@@
$\Phi(u)=\sum_{\lambda\in{\cal P}}{|\lambda\rangle}_{u}\widetilde{\otimes}\Phi_{\lambda}(u),$ (4.20)
$\Phi_{\lambda}(u)=\int_{\cal C}\prod_{a\in\lambda}dx_{a}:\prod_{a\in\lambda}x^{-}(x_{a}):\Phi_{\emptyset}(u)\prod_{\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\ \widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(\lambda;\mathfrak{z}),{}$
@@
$\Psi^{*}_{\lambda}(u)\xi=\Psi^{*}(u)(\xi\widetilde{\otimes}{|\lambda\rangle}_{u}),\qquad\forall\xi\in{\cal F}^{(1,N)}_{v},$ (4.21)
$\Psi^{*}_{\lambda}(u)=\int_{{\cal C}^{*}}\prod_{a\in\lambda}dx_{a}\ \widehat{\mathrm{Stab}}^{*}_{\mathfrak{C},T^{1/2}_{opp}}(\lambda;\mathfrak{z}^{*{-1}})\ :\prod_{a\in\lambda}x^{+}(x_{a}):\Psi^{*}_{\emptyset}(u)\prod_{\rho_{a}<\rho_{b}}<x^{+}(x_{a})x^{+}(x_{b})>^{sym}.{}$
The integration cycles ${\cal C},{\cal C}^{*}$ are chosen appropriately depending on the situation of the application. We call $\Phi_{\lambda}(u)$ ( resp. $\Psi^{*}_{\lambda}(u)$ ) with $\lambda\in{\cal P}_{n}$ the type I ( resp. type II dual ) vertex operator for $\mathrm{Hilb}^{n}({\mathbb{C}}^{2})$. The total operators $\Phi(u)$ and $\Psi^{*}(u)$ are then the vertex operators for $\bigoplus_{n}\mathrm{Hilb}^{n}({\mathbb{C}}^{2})$.
@@
$\Phi_{\lambda}(u)=\int_{\cal C}\prod_{a\in\lambda}dx_{a}\prod_{\rho_{a}\leq 0}x^{-}(x_{a})\cdot\Phi_{\emptyset}(u)\cdot\prod_{\rho_{a}>0}x^{-}(x_{a})$
×$\qquad\qquad\times\prod_{c_{a}=c_{b}\atop h_{a}>h_{b}}\frac{\theta(t_{2}x_{b}/x_{a})}{\theta(t_{1}x_{a}/x_{b})}\prod_{c_{a}+1=c_{b}\atop h_{a}<h_{b}}\frac{\theta(t_{2}x_{b}/x_{a})}{\theta(t_{1}x_{a}/x_{b})}\sum_{\delta\in\Upsilon_{\lambda}}W^{Ell}_{T^{1/2}}({\bf t}_{\delta};x_{1},\cdots,x_{|\lambda|},u;\mathfrak{z}),$
$\Psi^{*}_{\lambda}(u)=\int_{{\cal C}^{*}}\prod_{a\in\lambda}dx_{a}\ \sum_{\delta\in\Upsilon_{\lambda}}W^{Ell*}_{T^{1/2}_{opp}}({\bf t}_{\delta};x_{1},\cdots,x_{n},u;\mathfrak{z}^{*-1})\times\prod^{\curvearrowleft}_{\rho_{a}>0}x^{+}(x_{a})\cdot\Psi^{*}_{\emptyset}(u)\cdot\prod^{\curvearrowleft}_{\rho_{a}\leq 0}x^{+}(x_{a})$
×$\qquad\qquad\times\prod_{c_{a}=c_{b}\atop h_{a}>h_{b}}\frac{\theta^{*}(t_{2}x_{b}/x_{a})}{\theta^{*}(t_{1}x_{a}/x_{b})}\prod_{c_{a}+1=c_{b}\atop h_{a}<h_{b}}\frac{\theta^{*}(t_{2}x_{b}/x_{a})}{\theta^{*}(t_{1}x_{a}/x_{b})}.$
Here ${\displaystyle\prod_{a\in\lambda}x^{-}(x_{a})}$ (resp. ${\displaystyle\prod^{\curvearrowleft}_{a\in\lambda}x^{+}(x_{a})}$ ) denotes a product in the increasing (resp. decreasing ) order of $a\in\lambda$ from left to right.
@@
Note that the normal ordered products ${\displaystyle:\prod_{a\in\lambda}x^{\pm}(x_{a}):}$ are symmetric in $x_{a}\ a\in\lambda$. ∎
@@
4.3 Shuffle product formula
@@
4.3.1 Type I vertex operator
@@
$=\int_{{\cal C}}\prod_{a\in\lambda^{\prime}}dx^{\prime}_{a}\int_{{\cal C}}\prod_{b\in\lambda^{\prime\prime}}dx^{\prime\prime}_{b}:\prod_{a\in\lambda^{\prime}}x^{-}(x^{\prime}_{a}):\Phi_{\emptyset}(u_{1})\prod_{\rho_{a}<\rho_{b}}<x^{-}(x^{\prime}_{a})x^{-}(x^{\prime}_{b})>^{sym}\widehat{\mathrm{Stab}}_{\mathfrak{C}^{\prime},T^{1/2}}(\lambda^{\prime};\mathfrak{z})$
×$\qquad\times:\prod_{b\in\lambda^{\prime\prime}}x^{-}(x^{\prime\prime}_{b}):\Phi_{\emptyset}(u_{2})\prod_{\rho_{c}<\rho_{d}}<x^{-}(x^{\prime\prime}_{c})x^{-}(x^{\prime\prime}_{d})>^{sym}\widehat{\mathrm{Stab}}_{\mathfrak{C}^{\prime\prime},T^{1/2}}(\lambda^{\prime\prime};\mathfrak{z}).$
Here the chambers $\mathfrak{C}^{\prime},\mathfrak{C}^{\prime\prime}$ are the same and taken as the stability condition $t_{1}/t_{2}>0$. We also assume $|u_{1}|\ll|u_{2}|$. In a similar way to the type $A$ linear quiver case studied in [28], let us arrange the order of the elements in the integrand as follows.
@@
    Move $:\prod_{b\in\lambda^{\prime\prime}}x^{-}(x^{\prime\prime}_{b}):$ to the left of $\Phi_{\emptyset}(u_{1})$ by using the formula (4.16).
@@
    Make $:\prod_{a\in\lambda^{\prime}}x^{-}(x^{\prime}_{a})::\prod_{b\in\lambda^{\prime\prime}}x^{-}(x^{\prime\prime}_{b}):$ totally normal ordered product by the formula
    $:\prod_{a\in\lambda^{\prime}}x^{-}(x^{\prime}_{a})::\prod_{b\in\lambda^{\prime\prime}}x^{-}(x^{\prime\prime}_{b}):=\prod_{a\in\lambda^{\prime},b\in\lambda^{\prime\prime}}<x^{-}(x^{\prime}_{a})x^{-}(x^{\prime\prime}_{b})>\ :\prod_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}x^{-}(x_{a}):.$
@@
    Divide $<x^{-}(x^{\prime}_{a})x^{-}(x^{\prime\prime}_{b})>$ into the symmetric part (4.4) and the non-symmetric part (4.6).
@@
$\Phi_{\lambda^{\prime}}(u_{1})\Phi_{\lambda^{\prime\prime}}(u_{2})$ $=$ $\int_{{\cal C}\times{\cal C}}\prod_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}dx_{a}\ :\prod_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}x^{-}(x_{a}):\Phi_{\emptyset}(u_{1})\Phi_{\emptyset}(u_{2}){}$ (4.22)
×$\ \times\prod_{a,b\in(\lambda^{\prime},\lambda^{\prime\prime})\atop\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\ \widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}((\lambda^{\prime},\lambda^{\prime\prime});\mathfrak{z}),$

### v2

@@ 4.1. OPEs of the elliptic currents
$<x^{-}(u)x^{-}(v)>=\frac{(v/u,\hbar v/u,pv/t_{1}u,pv/t_{2}u;p)_{\infty}}{(t_{2}v/u,t_{1}v/u,pv/\hbar u,pv/u;p)_{\infty}},\quad|v/u|<1.$
@@
$<x^{+}(u)x^{+}(v)>^{sym}=t_{1}^{-1}\frac{(p^{*}t_{2}v/u,p^{*}t_{2}u/v,v/u,u/v;p^{*})_{\infty}}{(p^{*}\hbar v/u,p^{*}\hbar u/v,v/t_{1}u,u/t_{1}v;p^{*})_{\infty}},{}$
$\hskip 5.69054pt<x^{-}(u)x^{-}(v)>^{sym}=t_{1}^{-1}\frac{(pt_{1}^{-1}v/u,pt_{1}^{-1}u/v,\hbar v/u,\hbar u/v;p)_{\infty}}{(pv/u,pu/v,t_{2}v/u,t_{2}u/v;p)_{\infty}},{}$
@@
$<x^{+}(u)x^{+}(v)>^{sym}_{0}=\frac{(v/u,u/v,v/\hbar u,u/\hbar v;p^{*})_{\infty}}{(v/t_{1}u,u/t_{1}v,v/t_{2}u,u/t_{2}v;p^{*})_{\infty}},$
$<x^{-}(u)x^{-}(v)>^{sym}_{0}=\frac{(v/u,u/v,\hbar v/u,\hbar u/v;p)_{\infty}}{(t_{1}v/u,t_{1}u/v,t_{2}v/u,t_{2}u/v;p)_{\infty}},$
$<x^{+}(u)x^{+}(v)>^{non-sym}_{0}=\frac{\theta^{*}(t_{2}v/u)\theta^{*}(t_{1}v/u)}{\theta^{*}(v/u)\theta^{*}(\hbar v/u)},$
$<x^{-}(u)x^{-}(v)>^{non-sym}_{0}=\frac{\theta(t_{1}u/v)\theta(t_{2}u/v)}{\theta(u/v)\theta(\hbar u/v)}.$
@@
4.2. Vertex operators for ${\mathcal{M}}(n,1)$
We start from the vertex operators for ${\mathcal{M}}(n,1)$ i.e. $\mathrm{Hilb}^{n}({\mathbb{C}}^{2})$. Let us define
$\Phi_{\emptyset}(u):=(-u)^{-\alpha}e^{\Lambda_{0}}\;\widetilde{\Phi}_{\emptyset}(u),$ (4.7)
$\Psi^{*}_{\emptyset}(u):=(-u)^{\alpha}e^{-\Lambda_{0}-Q}\;\widetilde{\Psi}^{*}_{\emptyset}(u),$ (4.8)
@@ 4.2. Vertex operators for ${\mathcal{M}}(n,1)$
$\widetilde{\Phi}_{\emptyset}(u)=\exp\left\{-\sum_{m>0}\frac{1}{\kappa_{m}}\alpha^{\prime}_{-m}(\hbar^{1/2}u)^{m}\right\}$ (4.9)
$\hskip 56.9055pt\times\exp\left\{\sum_{m>0}\frac{1}{\kappa_{m}}\alpha^{\prime}_{m}(\hbar^{1/2}u)^{-m}\right\},{}$
$\widetilde{\Psi}^{*}_{\emptyset}(u)=\exp\left\{\sum_{m>0}\frac{1}{\kappa_{m}}\alpha_{-m}(\hbar^{1/2}u)^{m}\right\}$ (4.10)
$\hskip 56.9055pt\times\exp\left\{-\sum_{m>0}\frac{1}{\kappa_{m}}\alpha_{m}(\hbar^{1/2}u)^{-m}\right\}.{}$
@@
$\hskip 8.5359pt\Psi^{*}_{\emptyset}(u)x^{+}(v)=-\frac{\theta^{*}(v/u)}{\theta^{*}(\hbar u/v)}x^{+}(v)\Psi^{*}_{\emptyset}(u),$ (4.11)
$\hskip 8.5359pt\Phi_{\emptyset}(u)x^{-}(v)=-\frac{\theta(\hbar u/v)}{\theta(v/u)}x^{-}(v)\Phi_{\emptyset}(u),$ (4.12)
$\hskip 8.5359ptx^{+}(v)\Phi_{\emptyset}(u)=\Phi_{\emptyset}(u)x^{+}(v),\quad x^{-}(v)\Psi^{*}_{\emptyset}(u)=\Psi^{*}_{\emptyset}(u)x^{-}(v),$ (4.13)
$\hskip 8.5359pt\mathfrak{z}\Phi_{\emptyset}(u)=\hbar^{-1}\Phi_{\emptyset}(u)\mathfrak{z},\quad\Psi^{*}_{\emptyset}(u)\mathfrak{z}^{*}=\hbar\;\mathfrak{z}^{*}\Psi^{*}_{\emptyset}(u),$ (4.14)
$\hskip 8.5359pt[x^{\pm}(x),\mathfrak{z}]=[\Psi^{*}_{\emptyset},\mathfrak{z}]=0,\quad[x^{\pm}(x),\mathfrak{z}^{*}]=[\Phi_{\emptyset}(u),\mathfrak{z}^{*}]=0.$ (4.15)
@@
$\Phi(u)\ :\ {\mathcal{F}}^{(1,N)}_{v}\ \to\ {\mathcal{F}}^{(0,-1)}_{u}\widetilde{\otimes}{\mathcal{F}}^{(1,N+1)}_{-v/u}$,
$\Psi^{*}(u)\ :\ {\mathcal{F}}^{(1,N)}_{v}\widetilde{\otimes}{\mathcal{F}}^{(0,-1)}_{u}\ \to\ {\mathcal{F}}^{(1,N-1)}_{-vu}$
@@
$\Phi(u)=\sum_{\lambda\in{\mathcal{P}}}{|\lambda\rangle}_{u}\widetilde{\otimes}\Phi_{\lambda}(u),$ (4.16)
$\Phi_{\lambda}(u)=\int_{\mathcal{C}}\prod_{a\in\lambda}dx_{a}:\prod_{a\in\lambda}x^{-}(x_{a}):\Phi_{\emptyset}(u){}$
$\qquad\qquad\times\prod_{\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(\lambda;\mathfrak{z}){}$
@@
$\hskip 5.69054pt\Psi^{*}_{\lambda}(u)\xi=\Psi^{*}(u)(\xi\widetilde{\otimes}{|\lambda\rangle}_{u}),\qquad\forall\xi\in{\mathcal{F}}^{(1,N)}_{v}$, (4.17)
$\Psi^{*}_{\lambda}(u)=\int_{{\mathcal{C}}^{*}}\prod_{a\in\lambda}dx_{a}\ \widehat{\mathrm{Stab}}^{*}_{\mathfrak{C},T^{1/2}_{opp}}(\lambda;\mathfrak{z}^{*{-1}}):\prod_{a\in\lambda}x^{+}(x_{a}):\Psi^{*}_{\emptyset}(u){}$
$\qquad\qquad\times\prod_{\rho_{a}>\rho_{b}}<x^{+}(x_{a})x^{+}(x_{b})>^{sym}.{}$
The integration cycles ${\mathcal{C}},{\mathcal{C}}^{*}$ are chosen appropriately depending on the situation of the application. We call $\Phi_{\lambda}(u)$ ( resp. $\Psi^{*}_{\lambda}(u)$ ) with $\lambda\in{\mathcal{P}}_{n}$ the type I ( resp. the type II dual ) vertex operator for $\mathrm{Hilb}^{n}({\mathbb{C}}^{2})$. The total operators $\Phi(u)$ and $\Psi^{*}(u)$ are then the vertex operators for $\bigoplus_{n}\mathrm{Hilb}^{n}({\mathbb{C}}^{2})$.
One can take a cycle $\mathcal{C}$} (resp. $\mathcal{C}$$^{*}$ ) as an algebraic torus ${\mathbb{T}}^{|\lambda|}=\{(x_{a})\in{\mathbb{C}}^{|\lambda|}\ |\ |x_{a}|=1,\ a\in\lambda\ \}$, which encircles a half of the infinite series of poles in the OPE coefficients $x^{-}(x_{a})\Phi_{\emptyset}(u)>$ and $x^{-}(x_{a})x^{-}(x_{b})>$, $a,b\in\lambda$ satisfying $|x_{a}|>|x_{b}|$ for $\rho_{a}<\rho_{b}$ and $|x_{a}|>|u|$, $\forall a\in\lambda$ (resp. $x^{+}(x_{a})\Psi^{*}_{\emptyset}(u)>$ and $x^{+}(x_{a})x^{+}(x_{b})>$, $a,b\in\lambda$ satisfying $|x_{a}|>|x_{b}|$ for $\rho_{a}>\rho_{b}$ and $|x_{a}|>|u|$, $\forall a\in\lambda$). In Sec.5, we consider the $\mathrm{K}$-theoretic vertex functions and take the Jackson integral for $\int_{{\mathcal{C}}}$ and $\int_{{\mathcal{C}}^{*}}$. The elliptic stable envelopes $\widehat{\mathrm{Stab}}_{T^{1/2},\mathfrak{C}}(\bullet;\mathfrak{z})$ play a role of the pole subtraction matrices[1] and provide a basis of the cycles.
@@
$\Phi_{\lambda}(u)=\int_{\mathcal{C}}\prod_{a\in\lambda}dx_{a}\prod_{\rho_{a}\leq 0}x^{-}(x_{a})\cdot\Phi_{\emptyset}(u)\cdot\prod_{\rho_{a}>0}x^{-}(x_{a})$
$\qquad\qquad\times\prod_{c_{a}=c_{b}\atop h_{a}>h_{b}}\frac{\theta(t_{2}x_{b}/x_{a})}{\theta(t_{1}x_{a}/x_{b})}\prod_{c_{a}+1=c_{b}\atop h_{a}<h_{b}}\frac{\theta(t_{2}x_{b}/x_{a})}{\theta(t_{1}x_{a}/x_{b})}{}$
$\qquad\qquad\times\sum_{\delta\in\Upsilon_{\lambda}}W^{Ell}_{T^{1/2}}({\bf t}_{\delta};x_{1},\cdots,x_{|\lambda|},u;\mathfrak{z}),$
$\Psi^{*}_{\lambda}(u)=\int_{{\mathcal{C}}^{*}}\prod_{a\in\lambda}dx_{a}\ \sum_{\delta\in\Upsilon_{\lambda}}W^{Ell*}_{T^{1/2}_{opp}}({\bf t}_{\delta};x_{1},\cdots,x_{n},u;\mathfrak{z}^{*-1})$
$\qquad\qquad\times\prod^{\curvearrowleft}_{\rho_{a}>0}x^{+}(x_{a})\cdot\Psi^{*}_{\emptyset}(u)\cdot\prod^{\curvearrowleft}_{\rho_{a}\leq 0}x^{+}(x_{a})$
$\qquad\qquad\times\prod_{c_{a}=c_{b}\atop h_{a}>h_{b}}\frac{\theta^{*}(t_{2}x_{b}/x_{a})}{\theta^{*}(t_{1}x_{a}/x_{b})}\prod_{c_{a}+1=c_{b}\atop h_{a}<h_{b}}\frac{\theta^{*}(t_{2}x_{b}/x_{a})}{\theta^{*}(t_{1}x_{a}/x_{b})}.$
Here $\prod_{a\in\lambda}x^{-}(x_{a})$} (resp. $\prod^{\curvearrowleft}_{a\in\lambda}x^{+}(x_{a})$} ) denotes a product in the increasing (resp. decreasing ) order of $a\in\lambda$ from left to right.
@@
Note that the normal ordered products $:\prod_{a\in\lambda}x^{\pm}(x_{a}):$} are symmetric in $x_{a}\ a\in\lambda$. ∎
@@
4.3. Shuffle product and vertex operators for ${\mathcal{M}}(n,r)$
@@
4.3.1. The type I vertex operator
@@
$=\int_{{\mathcal{C}}}\prod_{a\in\lambda^{\prime}}dx^{\prime}_{a}\int_{{\mathcal{C}}}\prod_{b\in\lambda^{\prime\prime}}dx^{\prime\prime}_{b}:\prod_{a\in\lambda^{\prime}}x^{-}(x^{\prime}_{a}):\Phi_{\emptyset}(u_{1}){}$
$\qquad\times\prod_{\rho_{a}<\rho_{b}}<x^{-}(x^{\prime}_{a})x^{-}(x^{\prime}_{b})>^{sym}\widehat{\mathrm{Stab}}_{\mathfrak{C}^{\prime},T^{1/2}}(\lambda^{\prime};\mathfrak{z})$
$\qquad\times:\prod_{b\in\lambda^{\prime\prime}}x^{-}(x^{\prime\prime}_{b}):\Phi_{\emptyset}(u_{2})\prod_{\rho_{c}<\rho_{d}}<x^{-}(x^{\prime\prime}_{c})x^{-}(x^{\prime\prime}_{d})>^{sym}$
$\qquad\times\widehat{\mathrm{Stab}}_{\mathfrak{C}^{\prime\prime},T^{1/2}}(\lambda^{\prime\prime};\mathfrak{z}).$
Here the chambers $\mathfrak{C}^{\prime},\mathfrak{C}^{\prime\prime}$ are the same and taken as the stability condition $t_{1}/t_{2}>0$. We also assume $|u_{1}|>|u_{2}|$. In a similar way to the type $A$ linear quiver case studied in [28], let us arrange the order of the elements in the integrand as follows.
@@
    Move $\prod_{b\in\lambda^{\prime\prime}}x^{-}(x^{\prime\prime}_{b}):$ to the left of $\Phi_{\emptyset}(u_{1})$ by using the formula (4.12).
@@
    Make $\prod_{a\in\lambda^{\prime}}x^{-}(x^{\prime}_{a})::\prod_{b\in\lambda^{\prime\prime}}x^{-}(x^{\prime\prime}_{b}):$ totally normal ordered product by the formula
    $:\prod_{a\in\lambda^{\prime}}x^{-}(x^{\prime}_{a})::\prod_{b\in\lambda^{\prime\prime}}x^{-}(x^{\prime\prime}_{b}):$
    $=\prod_{a\in\lambda^{\prime},b\in\lambda^{\prime\prime}}<x^{-}(x^{\prime}_{a})x^{-}(x^{\prime\prime}_{b})>\ :\prod_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}x^{-}(x_{a}):.$
@@
    Divide $x^{-}(x^{\prime}_{a})x^{-}(x^{\prime\prime}_{b})>$ into the symmetric part (4.4) and the non-symmetric part (4.6).
@@
$\Phi_{\lambda^{\prime}}(u_{1})\Phi_{\lambda^{\prime\prime}}(u_{2})$ (4.18)
$=\int_{{\mathcal{C}}\times{\mathcal{C}}}\prod_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}dx_{a}\ :\prod_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}x^{-}(x_{a}):\Phi_{\emptyset}(u_{1})\Phi_{\emptyset}(u_{2}){}$
$\ \times\prod_{a,b\in(\lambda^{\prime},\lambda^{\prime\prime})\atop\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\ \widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}((\lambda^{\prime},\lambda^{\prime\prime});\mathfrak{z}),{}$